Grade 9General Physics

a body travel 200cm in the first 2sec$220cm in the 5sec.calculate the velocity at the end of the seventh sec from the start

Rahul
10 Years agoGrade 9

1 Answer

Vikas TU
10 Years ago
For first 2 sec.
200 = 2u + 0.5a*4 …...................(1)
220+200 = 7u + 0.5a*49
or
420 = 7u + 0.5a*49
or
60 = u + 3.5a …......................(2)
solve the 2 eqns. 
calculate u and a
and then apply
v = u + at for 7 seconds . i .e . t = 7s